# Linking Vision and Multi-Agent Communication through Visible Light Communication using Event Cameras

### 08 Feb 2024 v1

@@ 4.1. Settings
Based on the analysis in Section 3, it is plausible that there are scenarios where individual identification using RGB cameras is unattainable. As a result, we proceeded with the assumption that individual identification would not be feasible in the given scenario. Also, the location information was set to be unobtainable. In the present study, we have established the conditions in a two-dimensional plane using the Multi-Agent Particle Environments (MAPE) Lowe et al. (2017) environment, where one-dimensional visual information can be obtained (Appendix). In the VLC configuration, communication info was assumed to come only from visible agents. Obstructions by agents or landmarks prevent signal acquisition in this setup. The communication method used in the present study with event-VLC provides communication information linked to visual information. Therefore, in the case of the event-VLC, individual identification is assumed to be possible.
@@ 4.3. Models and Algorithms
Simple Spread, Predator-Prey First, for the tasks of Simple Spread, Predator-prey, we employed a reinforcement learning algorithm that incorporates limited vision Nakagawa and Kanezaki (2023). As illustrated in Fig. 5, agents learn visual direction changes separately from movement actions. In terms of learning actions, reinforcement learning is conducted through the MADDPG Lowe et al. (2017) algorithm to determine how much to advance or when to stop. The models utilized for learning behaviors involve a one-dimensional convolutional layer shared between the actor and critic networks, processing one-dimensional visual data (Appendix).
@@
Simple Swing For the Simple Swing task, which does not involve movement, we partially employed the methodology mentioned above and adapted it for experimental conditions. Each agent retains information about its visual direction and the entities it can see based on visual data. Within the critic model, individual observations and actions are combined, and network is trained to predict the next self topological information after altering the visual direction by actions. Details of the model are included in the Appendix. The loss function of predicting the next topological information is represented by the following equation:
Subsequently, the concatenated vector is processed by the MLP network to output Q values. Learning using TD error loss is backpropagated only through the network . Critic’s loss function in the reinforcement learning algorithm used in this study is described in Appendix. Actor is trained solely through MLP from their individual observations. In this task, simplified information about what is visible, rather than high-dimensional visual data, is utilized as observations. This approach aims to conduct physical experiments without considering the Sim2Real aspects.
@@ 4.4. Results
The results of 10K times benchmark on the trained model are shown in Fig. 6. The trajectory of the agent in tasks requiring movement is depicted in Fig of the Appendix. First, for the Simple Spread task, the performance of the Event-VLC system was superior. This can be attributed to the ability to comprehend the surroundings by linking which agent conveys communication information when coordinating limited visual information in the case of event-VLC. This trend was similarly observed in the Predator-prey task. The adversarial agent endeavors to escape toward the outer boundaries by maneuvering through the spaces between the good agents. However, the good agents collaborate visually, enabling them to pinpoint the position of the adversarial agent swiftly.
Next, concerning the Simple Swing task, the performance of Event-VLC was comparable to that of radio communication. Despite the disadvantageous setting for VLC, where a limited field of view makes communication with other agents difficult, the ability of event-VLC to associate location and visual information in communication compensates for this, resulting in comparable performance to radio communication, which can obtain all communication information. When training with the ability to predict the next visual input, as utilized in the model architecture in the present paper, the performance surpasses simple training through MLP. Furthermore, the configuration of providing rewards for VLC communication itself also contributes to the overall performance improvement. More details are provided in Appendix.
@@
Additionally, as indicated in Table in the Appendix, in an environment where agents can always obtain information about other agents through radio communication and be identified, the highest performance is achieved. This indicates that factors such as the inability to distinguish opponents and lack of constant access to communication information are elements that degrade performance. As Fig. 6 shows, the performance degradation due to the lack of individual identification and the inability to link with communication information is significant. In this chapter, it was demonstrated that depending on the task setup, performance can be enhanced with event-VLC approaches.
@@ 5.1. Settings
The multi-agent operation of one step is illustrated in Fig. 7. Agents determined which landmarks were within their field of view using RGB cameras. They encoded the visual direction and the visible landmark and agent into binary signals, transmitted as flashes of LED lights. Each agent decoded these light signals using event cameras and combined the received information with its own and other agents’ information from the past three frames. Using the trained network, agents selected their next visual direction. This system was designed using Robot Operating System (ROS), and the ROS node relationships are depicted in the Appendix.

### 15 Feb 2024 v2

@@ 4.1. Settings
Based on the analysis in Section 3, it is plausible that there are scenarios where individual identification using RGB cameras is unattainable. As a result, we proceeded with the assumption that individual identification would not be feasible in the given scenario. Also, the location information was set to be unobtainable. In the present study, we have established the conditions in a two-dimensional plane using the Multi-Agent Particle Environments (MAPE) Lowe et al. (2017) environment, where one-dimensional visual information can be obtained (Appendix A). In the VLC configuration, communication info was assumed to come only from visible agents. Obstructions by agents or landmarks prevent signal acquisition in this setup. The communication method used in the present study with event-VLC provides communication information linked to visual information. Therefore, in the case of the event-VLC, individual identification is assumed to be possible.
@@ 4.3. Models and Algorithms
Simple Spread, Predator-Prey First, for the tasks of Simple Spread, Predator-prey, we employed a reinforcement learning algorithm that incorporates limited vision Nakagawa and Kanezaki (2023). As illustrated in Fig. 5, agents learn visual direction changes separately from movement actions. In terms of learning actions, reinforcement learning is conducted through the MADDPG Lowe et al. (2017) algorithm to determine how much to advance or when to stop. The models utilized for learning behaviors involve a one-dimensional convolutional layer shared between the actor and critic networks, processing one-dimensional visual data (Appendix A).
@@
Simple Swing For the Simple Swing task, which does not involve movement, we partially employed the methodology mentioned above and adapted it for experimental conditions. Each agent retains information about its visual direction and the entities it can see based on visual data. Within the critic model, individual observations and actions are combined, and network is trained to predict the next self topological information after altering the visual direction by actions. Details of the model are included in the Appendix B. The loss function of predicting the next topological information is represented by the following equation:
Subsequently, the concatenated vector is processed by the MLP network to output Q values. Learning using TD error loss is backpropagated only through the network . Critic’s loss function in the reinforcement learning algorithm used in this study is described in Appendix C. Actor is trained solely through MLP from their individual observations. In this task, simplified information about what is visible, rather than high-dimensional visual data, is utilized as observations. This approach aims to conduct physical experiments without considering the Sim2Real aspects.
@@ 4.4. Results
The results of 10K times benchmark on the trained model are shown in Fig. 6. The trajectory of the agent in tasks requiring movement is depicted in Fig. 12 of the Appendix E. First, for the Simple Spread task, the performance of the Event-VLC system was superior. This can be attributed to the ability to comprehend the surroundings by linking which agent conveys communication information when coordinating limited visual information in the case of event-VLC. This trend was similarly observed in the Predator-prey task. The adversarial agent endeavors to escape toward the outer boundaries by maneuvering through the spaces between the good agents. However, the good agents collaborate visually, enabling them to pinpoint the position of the adversarial agent swiftly.
Next, concerning the Simple Swing task, the performance of Event-VLC was comparable to that of radio communication. Despite the disadvantageous setting for VLC, where a limited field of view makes communication with other agents difficult, the ability of event-VLC to associate location and visual information in communication compensates for this, resulting in comparable performance to radio communication, which can obtain all communication information. When training with the ability to predict the next visual input, as utilized in the model architecture in the present paper, the performance surpasses simple training through MLP. Furthermore, the configuration of providing rewards for VLC communication itself also contributes to the overall performance improvement. More details are provided in Appendix D.
@@
Additionally, as indicated in Table 4 in the Appendix E, in an environment where agents can always obtain information about other agents through radio communication and be identified, the highest performance is achieved. This indicates that factors such as the inability to distinguish opponents and lack of constant access to communication information are elements that degrade performance. As Fig. 6 shows, the performance degradation due to the lack of individual identification and the inability to link with communication information is significant. In this chapter, it was demonstrated that depending on the task setup, performance can be enhanced with event-VLC approaches.
@@ 5.1. Settings
The multi-agent operation of one step is illustrated in Fig. 7. Agents determined which landmarks were within their field of view using RGB cameras. They encoded the visual direction and the visible landmark and agent into binary signals, transmitted as flashes of LED lights. Each agent decoded these light signals using event cameras and combined the received information with its own and other agents’ information from the past three frames. Using the trained network, agents selected their next visual direction. This system was designed using Robot Operating System (ROS), and the ROS node relationships are depicted in the Appendix F.
@@ Acknowledgements
Appendices
A. The One-Dimensional Visual Information and Basic Processing Network
In the simulation, we utilized one-dimensional visual information Fig. 8. In the event-VLC approach, it is assumed that communication signals via VLC can be obtained in the same field of view as RGB information. In cases where the Field of View (FOV) is less than or equal to 120, the parameter is set to 90. Conversely, when the FOV is 360, the parameter is set to 360. The field size was set as a square measuring 80 cm on each side, with agents and landmarks sized at 6 cm.
The basic learning model uses the Actor-Critic method. In the critic component, centralized training is performed using the one-dimensional visual information from all agents. The individual one-dimensional visual information is combined through a one-dimensional convolutional network to output Q-values. As for the actor component, decentralized execution is carried out using individual information and communication data only. The one-dimensional convolutional network of actor and critic is composed of four-layer 1D-convolutional with 64 filters and a kernel size of 4. The MLP layers are composed of a five-layer MLP with 256 units. In addition, the and networks are constructed with four-layer MLP with 256 units.
B. Model for Simple Swing
For Simple Swing task, which does not involve movement and involve coordination of visual directions, we introduced a network model into the critic that predicts what will be seen when the visual direction is changed. In our approach, the network predicts the next topological information and combines its output to compute Q values in the network . During the training process, the network is trained using the prediction loss of the next topological information, while the network is trained using the TD loss as its loss function.
C. Basic Reinforcement Learning Algorithm
In this study, we consider the behavior of multi-agent as a Markov decision process (MDP). For a game of agents, we describe the state of all agents as . We also denote the observed value of each agent as , and each agent takes an action .
Letting be the measure for the agent to decide on an action, is a parameter of the neural network. This policy is defined by . Each agent’s action acts on the environment and moves on to the next . This is the state transition function when there are agents’ actions. Each agent is then given a reward when it takes action in state . The discount sum is defined as follows using the discount rate .
One method of reinforcement learning is the policy gradient method. These optimize the agent’s parameters so as to maximize the objective function . The gradient of this objective function can be derived as follows.
Q value in the equation can be calculated by various methods. In this paper, we use the Actor-Critic algorithm, in which the critic learns to predict by TD error. Also, the measure gradient method can be extended to deterministic algorithms. In deterministic policies, the action for a state is determined according to . When deterministic action is adopted, the gradient of the objective function can be derived as follows.
In this paper, deterministic policies are employed and the basic learning algorithm is same as MADDPG.
Task Name |
event-VLC | 86 ± 0.45 [%] | 8.0 ± 0.34 [step] | 97 ± 0.42 [%] | 0.097 ± 0.0016 [m] | 98 ± 0.41 [%] | ||||||||||
radio | 83 ± 0.50 [%] | 9.5 ± 0.42 [step] | 98 ± 0.33 [%] | 0.14 ± 0.0021 [m] | 97 ± 0.50 [%] | ||||||||||
w/o comm. | 81 ± 0.50 [%] | 12 ± 0.38 [step] | 93 ± 0.63 [%] | 0.20 ± 0.0025 [m] | 87 ± 0.85 [%] | ||||||||||
radio (w/Visual ID) | 88 ± 0.58 [%] | 7.8 ± 0.26 [step] | 99 ± 0.33 [%] | 0.098 ± 0.0020 [m] | 98 ± 0.39 [%] |
D. Ablation Study for Simple Swing
We present an ablation study of the reward setting for the VLC itself and a predictive network model for the next topological information introduced for the Simple Swing task. Thus, we can see the impact of increasing overall performance. From these results, it can be inferred that providing rewards for performing VLC itself and employing a network structure that predicts the next topological information contribute to performance improvement. Providing rewards for VLC itself enables learning of policies that facilitate moderate communication and prevent using only one’s own information.
E. Quantitative benchmark results
The results of Fig. 6 are summarized quantitatively in Table 4. As can be seen, the VLC method performs better in settings where the appearance of other agents is indistinguishable than in the case of radio communication or no communication. On the other hand, it performs best when radio communication plus other agents are distinguishable. Furthermore, Fig. 12 illustrating the trajectories of agents for tasks involving movement in Simple Spread, Predator-Prey, Target Encirclement and Goal Crossing. Fig. 12 shows the behavior of the agents trained in this simulation for tasks involving movement. In each task, the agents are able to cooperate with each other to accomplish the task.
F. The Relationship of ROS Nodes in Multi-Agent Experiments
Fig. 13 illustrates the relationships of each ROS node in three agents. RGB camera and Event camera acquire information through usb_cam and event_publisher, respectively. The images captured by the RGB camera are processed by the landmark node, which maintains information about visible landmarks. Regarding Event information, processed data from the decoder node and the angle of the mobility node’s turntable are sent to the led_communication_node, which emits signals from LEDs. Similarly, the data is sent to the integration_msg_node to calculate the next rotation angle in the policypi node. Certain nodes contained specific terms related to the experiment, which have been revised in red to generic terms.